OpenStudy (anonymous):

The Next Step in Solving: 5x - 30 + 6x = 18x +54 Please?

OpenStudy (texaschic101):

I suppose the next step is combining like terms 5x - 30 + 6x = 18x + 54 -- combine like terms 11x - 30 = 18x + 54 -- subtract 18x from both sides 11x - 18x - 30 = 54 -- add 30 to both sides 11x - 18x = 54 + 30 -- combine like terms -7x = 84 -- divide by -7 x = - 12 check.. 11x - 30 = 18x + 54 11(-12) - 30 = 18(-12) + 54 -132 - 30 = -216 + 54 -162 = - 162 (correct)

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

anytime :)